OpenStudy (anonymous):

I need help with Laplace Transforms. Solve the system: x'+2y'+x=0 x'-y'+y=0 x(0)=0 y(0)=1 I got to the part where sX+2sY+X-2=0 and sX-sY+Y+1=0 How do I solve the system from here? thanks!

OpenStudy (mr.math):

You have \(sX+2sY+X-2=0\) and \(sX-sY+Y+1=0\)

OpenStudy (mr.math):

Solve them as you do with any other system of equation, either by substitution or elimination.

OpenStudy (mr.math):

Your variables here are X and Y.

OpenStudy (anonymous):

\[\left\{X=-\frac{2}{3 s^2-1},Y=-\frac{-3 s-1}{3 s^2-1}\right\} \]